In an effort to stimulate the economy, the Federal Reserve decides to buy $5 million worth of gold. Assuming the required reserve rate is 14%, what is the maximum amount of total money that could result from this initial injection?

Question 10 options:

a)

$4,250,000

b)

$5,000,000

c)

$9,250,000

d)

$33,333,333.33

e)

None of the above.
Answer #1

The multiplier in this economy = 1/r = 1/0.14 = 7.1428

The Fed buy $5 million worth of gold

so, the maximum amount of money supply that could result from this would be:-

= 5 million*7.1428

= 35.71 million

option(E)